# Biomedical Event Extraction via Structure-aware Generation

### v1

@@ IV Experiments
This section first describes the dataset, evaluation metrics, implementation details, and baseline models. Then, we present the experimental results to show the effectiveness of the proposed GenBEE model for biomedical event extraction.
@@ IV-A Datasets
We have conducted experiments on the following three publicly available benchmark datasets, namely MLEE, GE11, and PHEE, for biomedical event extraction. MLEE [19] has 29 event types and 14 argument roles. We use the train/dev/test split given by the data provider. GE11 [20] has 9 event types and 10 role arguments. We use the train/dev/test split given by the shared task and evaluate the performance based on the development set, as the test set is unannotated and the official tool for evaluation is no longer available. PHEE [30] has 2 event types and 16 role arguments. We use the train/dev/test split provided by TextEE [31]. We preprocess each dataset by using the script of TextEE [31]. Table III presents the details of the splits and the datasets. For evaluation metrics, we follow the previous work [32, 10, 13] to adopt two metrics to evaluate performance. (1) Trigger Classification (Trg-C) F1-score: A trigger is correctly classified if the predicted span and the predicted event type match the gold ones. (2) Argument Classification (Arg-C) F1 score: An argument is correctly classified if the predicted span, event type, and role type match the gold ones.
@@ IV-C Baseline Models
TABLE IV: Experimental results (%) for extracting events based on the MLEE, GE11 and PHEE datasets.
@@ IV-D Experimental Results
Table IV shows the Trg-C F1-scores and Arg-C F1-scores based on the MLEE, GE11, and PHEE datasets. The best score is highlighted in bold, and the second-best score is underlined. Generation-based models and Classification-based models are indicated by “Gen” and “Cls”, respectively. The letter l in the column PLM denotes the large model. Overall, the GenBEE model achieves new state-of-the-art performance on the MLEE and GE11 datasets. Moreover, on the PHEE dataset, our model achieves the second-highest Arg-C F1-score among all baselines, while its Trg-C F1-score is still competitive with other classification-based models. Additionally, we reckon that the simpler event structures with less structural information in the PHEE datasets are the main reason that affects the performance of the GenBEE model.
Table IV also shows that GenBEE significantly outperforms the state-of-the-art generation-based model, DEGREE, in experiments based on the MLEE, GE11, and PHEE datasets. More specifically, GenBEE demonstrates an improvement of 3.4%, 3.0%, and 2.2% in Trg-C F1-score, and an improvement of 3.3%, 3.9% and 2.4% in Arg-C F1-score, over DEGREE on the MLEE, GE11, and PHEE datasets, respectively. We attribute GenBEE’s performance improvement to the use of event prompts and structural prefixes, which provide event structural information for biomedical event extraction.
For the experiments on the PHEE dataset, it is worth noting that the two state-of-the-art models, TagPrime and DyGIE++, are both sequence tagging models. We find that these two models mainly benefit from better span identification of trigger and argument words for achieving better performance. Even though the Trg-C F1-score of GenBEE slightly lags behind some baselines, GenBEE still achieves the second-highest Arg-C F1-score among all baselines. We attribute this to GenBEE’s end-to-end extraction style, which can be less susceptible to the error propagation issue that commonly affects pipeline methods.
TABLE V: Ablation study on the proposed GenBEE model based on different model configurations. We report numbers in F1-scores (%).
@@
(1) -w/o Structural Prompts 80.8 67.2 67.5 63.5 69.6 53.4 -0.5 -0.7
(2) -w/o Structural Prefix 79.8 66.7 66.4 62.1 68.7 52.6 -1.5 -1.6
(3) -w/o Event Prompts 80.9 67.3 67.5 63.0 69.2 52.9 -0.6 -1.0
(4) -w/o (1) (2) (3) 78.0 64.6 65.2 60.5 67.6 51.4 -2.9 -3.2
V Analysis
V-A Ablation Study
We conduct an ablation study of the proposed GenBEE model to evaluate the effects of different components on the overall performance across the MLEE, GE11, and PHEE benchmark datasets. Specifically, we focus on the following three components: structural prompts, structural prefixes, and event prompts. Table V reports the performance results of the ablation study. The symbol $\triangle$ indicates the difference in the F1-score between the different configuration models and the proposed GenBEE model. As shown in Table V, we observe that the structural prompts, structural prefixes, and event prompts significantly contribute to the performance of our proposed model, regardless of the dataset. Firstly, removing structural prompts results in a drop of 0.5% and 0.7% in the average Trg-C and Arg-C F1-scores, respectively. This demonstrates that structural prompts can boost the performance of the proposed model. Secondly, the removal of the structural prefix from GenBEE results in a reduction of 1.5% and 1.6% in the average Trg-C and Arg-C F1-scores, respectively, highlighting the important role of this component in the model’s overall performance. Removing the event prompts further leads to a decrease of 0.6% and 1.0% in the average Trg-C and Arg-C F1-scores, respectively. This drop in performance indicates that the prompts constructed for describing events significantly contribute to the improvement of the proposed GenBEE model. Overall, each component of the GenBEE model plays a crucial role in achieving promising performance for biomedical event extraction.
@@ VI Conclusion
We propose GenBEE, a novel generative model with structure-aware prefixes for biomedical event extraction. The experimental results demonstrate that GenBEE outperforms strong baselines on the MLEE, GE11, and PHEE datasets. Moreover, our experiments show that prefixes effectively serve as a medium to link structural prompts with the representation space of generative models, thereby enhancing the model’s overall performance.

### v2

@@ IV Experiments
In this section, we first describe the dataset, evaluation metrics, implementation details, and baseline models. Then, we present the experimental results to show the effectiveness of the proposed GenBEE model for biomedical event extraction.
@@ IV-A Datasets
We have conducted experiments on the following three publicly available benchmark datasets, namely MLEE, GE11, and PHEE, for biomedical event extraction. MLEE [19] has 29 event types and 14 argument roles. We use the train/dev/test split given by the data provider. GE11 [20] has 9 event types and 10 role arguments. We use the train/dev/test split given by the shared task and evaluate the performance based on the development set, as the test set is unannotated and the official tool for evaluation is no longer available. PHEE [30] has 2 event types and 16 role arguments. We use the train/dev/test split provided by TextEE [31]. We preprocess each dataset by using the script of TextEE [31]. Table III presents the statistics of the datasets. For evaluation metrics, we follow the previous work [32, 10, 13] to adopt the following two metrics to evaluate the performance: (1) Trigger Classification (Trg-C) F1-score: A trigger is correctly classified if the predicted span and the predicted event type match the gold ones. (2) Argument Classification (Arg-C) F1 score: An argument is correctly classified if the predicted span, event type, and role type match the gold ones.
@@ IV-C Baseline Models
TABLE IV: Experimental results (%) for extracting events based on the MLEE, GE11 and PHEE datasets. The best score is highlighted in bold, and the second-best score is underlined.
@@ IV-D Experimental Results
Table IV shows the Trg-C F1-scores and Arg-C F1-scores based on the MLEE, GE11, and PHEE datasets. “Gen” and “Cls” denote Generation-based models and Classification-based models, respectively. The letter “l” in the column PLM is used to denote the large model. Overall, the GenBEE model achieves new state-of-the-art performance on the MLEE and GE11 datasets. Moreover, on the PHEE dataset, our model achieves the second-highest Arg-C F1-score among all baselines, while its Trg-C F1-score is still competitive with other classification-based models. Additionally, we reckon that the simpler event structures with less structural information in the PHEE datasets are the main reason that affects the performance of the GenBEE model.
Table IV also shows that GenBEE significantly outperforms the state-of-the-art generation-based model, DEGREE, in the experiments based on the MLEE, GE11, and PHEE datasets. More specifically, GenBEE demonstrates an improvement of 3.4%, 3.0%, and 2.2% in Trg-C F1-score, and an improvement of 3.3%, 3.9% and 2.4% in Arg-C F1-score, over DEGREE on the MLEE, GE11, and PHEE datasets, respectively. We attribute GenBEE’s performance improvement to the use of event prompts and structural prefixes, which provide event structural information for biomedical event extraction.
For the experiments on the PHEE dataset, it is worth noting that the two state-of-the-art models, TagPrime and DyGIE++, are both sequence tagging models. We find that these two models mainly benefit from better span identification of trigger and argument words for achieving better performance. Even though the Trg-C F1-score of GenBEE slightly lags behind some baselines, GenBEE still achieves the second-highest Arg-C F1-score when compared with all the baselines. We attribute this to GenBEE’s end-to-end extraction style, which can be less susceptible to the error propagation issue that commonly affects pipeline methods.
V Analysis
V-A Ablation Study
We conduct an ablation study of the proposed GenBEE model to evaluate the effects of different components on the overall performance across the MLEE, GE11, and PHEE benchmark datasets. Specifically, we focus on the following two components: Structural Prefix Learning and Event Prompt Construction. Table V reports the performance results of the ablation study. The symbol $\triangle$ indicates the difference in the F1-score between the different configuration models and the proposed GenBEE model. As shown in Table V, we observe that the structural prefixes and event prompts significantly contribute to the performance of our proposed model, regardless of the dataset. Firstly, the removal of the Structural Prefix Learning module from GenBEE results in a reduction of 1.5% and 1.6% in the average Trg-C and Arg-C F1-scores, respectively, highlighting the important role of this component in the model’s overall performance. Moreover, removing the event prompts leads to a decrease of 0.6% and 1.0% in the average Trg-C and Arg-C F1-scores, respectively. This drop in performance indicates that the prompts constructed for describing events significantly contribute to the improvement of the proposed GenBEE model. Overall, each component of the GenBEE model plays a crucial role in achieving promising performance for biomedical event extraction.
TABLE V: Ablation study on the proposed GenBEE model based on different model configurations. We report the results in F1-scores (%).
@@ V-A Ablation Study
(1) -w/o Structural Prefix Learning 79.8 66.7 66.4 62.1 68.7 52.6 -1.5 -1.6
(2) -w/o Event Prompts Construction 80.9 67.3 67.5 63.0 69.2 52.9 -0.6 -1.0
(3) -w/o (1) (2) 78.0 64.6 65.2 60.5 67.6 51.4 -2.9 -3.2
@@ VI Conclusion
We propose GenBEE, a novel generative model with structure-aware prefixes for biomedical event extraction. The experimental results demonstrate that GenBEE outperforms the strong baselines on the MLEE, GE11, and PHEE datasets. Moreover, our experiments show that structure-aware prefixes can effectively serve as a medium to link structural prompts with the representation space of generative models, thereby enhancing the model’s overall performance.